# Machine-learning based high-bandwidth magnetic sensing

### v1

@@ Abstract
Recent years have seen significant growth of quantum technologies, and specifically quantum sensing, both in terms of the capabilities of advanced platforms and their applications. One of the leading platforms in this context is nitrogen-vacancy (NV) color centers in diamond, providing versatile, high-sensitivity, and high-resolution magnetic sensing. Nevertheless, current schemes for spin resonance magnetic sensing (as applied by NV quantum sensing) suffer from tradeoffs associated with sensitivity, dynamic range, and bandwidth. Here we address this issue, and implement machine learning tools to enhance NV magnetic sensing in terms of the sensitivity/bandwidth tradeoff in large dynamic range scenarios. We experimentally demonstrate this new approach, reaching an improvement in the relevant figure of merit by a factor of up to 5. Our results promote quantum machine learning protocols for sensing applications towards more feasible and efficient quantum technologies.
@@
In the context of quantum sensing, one of the leading systems is based on Nitrogen Vacancy (NV) color centers in diamond [1], which provides a versatile platform for diverse quantum sensing, notably magnetic sensing [2, 3, 4, 5, 6]. NVs have found important applications in magnetic sensing, covering paleomagnetometry [7], biosensing [8, 9, 10, 11], nuclear magnetic resonance [12, 13, 14] and more.
Quantum sensing with NVs is realized through spin resonance measurements, usually detected optically. While this approach achieves quantitative vectorial information with high sensitivity and spatial resolution, it suffers from a trade-off between sensitivity and bandwidth, specifically in the high dynamic range regime. In fact, working with small fields (small dynamic range) enables an optimal sensing strategy, which relies on precise measurements at a predetermined high-sensitivity point (the point of maximal signal gradient). However, this is not possible in the regime of large dynamic range signals.
In this paper, we address this limitation and introduce a Machine Learning (ML) algorithm that significantly improves this trade-off. We demonstrate that training an appropriate neural network using a combination of real and simulated data enables a clear improvement in measurement bandwidth for a given sensitivity goal, in the large dynamic range scenario.
Neural Networks are powerful ML models constituted by interconnected layers of artificial neurons that are trained to minimize a loss function [15, 16]. Such models have been already used for sensing of magnetic fields and for noise spectroscopy with NV centers [17, 18, 19]. By contrast, in this work neural networks find application in the full reconstruction of an external magnetic field. We analyze the network’s performance for a decreasing number of measurement points, while exploring different training dataset sizes, and altering noise and lineshape conditions. We compare these results to similar analysis done for raster scans and find that the network has advantages over raster scanning. With the right training, the network can be insensitive to variations in noise and lineshape. Moreover, the network’s measurement error is more robust and scales better for fewer data points.
As shown in Fig. 1, the standard approach to quantitatively measure vectorial magnetic fields using NVs is essentially a common spin resonance measurement termed Electron Spin Resonance (ESR), usually performed through optical readout for NVs, referred to as Optically Detected Magnetic Resonance (ODMR), using an experimental setup schematically depicted in Fig. 1a [3, 2].
@@
Figure 1: LABEL:sub@fig:scheme-setup Experimental setup schematic; MW pulses are used to find the spin resonances of the NV center within the diamond sample related to the external magnetic field orientation and intensity, green excitation laser illuminates the diamond and fluorescence from the NV centers is collected and read by a detector. LABEL:sub@fig:scheme-nv The NV can appear in four possible orientations within the diamond lattice. LABEL:sub@fig:scheme-level NV energy level schematic: triplet ground state; under an external magnetic field the degeneracy between the $m_{s}=\pm 1$ energy levels is lifted. MW drives transitions between $m_{s}=0$ and $m_{s}=\pm 1$. Excitation of the ground state with Green laser (532 nm) can decay in two ways: one is non-radiative and non-spin conserving through the singlet state. The second path is spin-conserving and radiates in red (650-800 nm). The $m_{s}=0$ state decays mostly radiatively, while $m_{s}=\pm 1$ includes a significant non-radiative decay through the singlet state into the ground spin state $m_{s}=0$. LABEL:sub@fig:scheme-esr ESR spectrum, with measured data in blue and the corresponding fit in purple (to a set of eight Lorentzian functions). In black, a subset of the data, i.e $\lambda_{i}$. The resonance frequencies are marked with $y_{i}$. LABEL:sub@fig:scheme-net Schematic of the neural network, the inputs $\lambda_{i}$ are the measured data points and the outputs are the predictions $y^{\prime}_{i}$ of the eight resonance frequencies $y_{i}$.
@@
The NV defect orientation within the crystal lattice (Fig. 1b) and its energy level structure (Fig. 1c), enable the optical detection of the NV spin resonances, realized through the combined application of green light excitation and microwave (MW) irradiation, while detecting the NV spin-dependent red fluorescence. As depicted in Fig. 1d, a full ESR spectrum of an NV ensemble consists of eight resonances, representing two resonant frequencies for each one of the four crystallographic orientations the NV can take in a single crystal diamond sample (Fig. 1b). Basically, this measurement amounts to identifying resonance positions (i.e. $y_{i}$) in the frequency space. Once obtained, a full vectorial magnetic field can be extracted. Further details regarding the NV system, the experimental setup (Figs. 1a and 1c) and the ODMR measurement scheme can be found in Sec. III.
@@
This process of identifying the resonance positions requires scanning the signal over frequency space (sometimes referred to as raster scanning). This scan should cover the entire relevant frequency range (determined by the desired dynamic range), and the scan resolution (number of data points within the frequency window) is determined by the resonance widths and Signal to Noise Ratio (SNR). There is a minimal scan resolution required to successfully retrieve the resonance frequencies, and in general the measurement error depends on these various parameters (linewidth, SNR, scan resolution), as detailed in Sec. III.
@@
The sensitivity for an ODMR measurement is defined to be $\eta=(\delta\nu*\sqrt{T})/\gamma$, where $\delta\nu$ is the error for the resonance frequency, T is the measurement time (essentially corresponding to the averaging time and affecting the SNR) and $\gamma$ is the gyromagnetic ratio of the NV [20]. A shorter measurement time both improves sensitivity and allows for faster measurements, which could be crucial for time varying signals. Reducing measurement time could be achieved by improving the experimental setup, i.e, having higher SNR or better NV properties. However, it is not less important to consider the optimal number of data points in a measurement, as it essentially scales linearly with measurement time.
ML is a research field that deals with the development of artificial intelligence methods that learn from data. Several approaches were developed to solve the two main supervised ML tasks of classification, i.e. to predict a categorical value associated to some predictors (called features in the ML jargon), and regression where the predicted value is continuous [21]. Supervised learning means that the model is trained to learn to predict based on a ground truth associated to each input. By contrast, in unsupervised learning the dataset is composed only by features without any desired output [21]. In the latter, the main task is called clustering where the model learns to divide the data into groups based on the elements distance. The third ML approach is called reinforcement learning [22] where the model is represented by an agent that learns to implement a policy to decide actions in a simulated environment. Some of the most common classical methods are K-Means for clustering, K-Nearest Neighbor (KNN), decision trees, Support Vector Machine (SVM) and logistic regression for classification tasks and linear regression for regression tasks [23, 21]. Neural Networks were inspired by early research on brain modeling and successively find application as a ML tool. Recently, with the increment of computational power and the increasing availability of large datasets, they overcome other ML algorithms as a general-purpose tool for either clustering, classification and regression tasks. In fact they are capable of universal function approximation [24] and easily adapt to different scenarios. Moreover, increasing the depth of the network allows the model to learn to extract higher level features from the input, and in this context they are denoted by deep learning [15, 16]. There are several neural network based models specialized for specific kinds of data, specifically Convolutional Neural Networks (CNN[25] for images and Recurrent Neural Networks (RNN[26] for time series and text. Remarkably, in the latter years we are witnessing the rising of generative models for images with Generative Adversarial Networks (GAN[27], diffusion models [28, 29, 30], and text with Large Language Models (LLM[31].
Quantum Machine Learning (QML) is a growing research field at the intersection of ML and quantum physics and computing [32, 33]. In general, it deals with different approaches depending on the nature of the analyzed data and of the models used to perform it. Classical ML models can be successfully adopted to process quantum data for instance, with the objective to analyze the noise affecting a quantum device [19, 34, 35] or to control quantum dynamics [36]. QML models can be also implemented on quantum devices to perform tasks on classical data [37] or even on fully quantum environments [38].
In particular, in this work we adopt a classical Multilayer Perceptrons in several supervised regression tasks on quantum data. MLPs are neural network models organized in a feed-forward structure [21, 16, 15] as in Fig. 1e. MLPs are trained via gradient descent to minimize a loss function between the prediction and the desired output. More details on the used models are given in Sec. III.2.
In the following, we detail our approach of training the ML models to identify the resonance frequencies, given a limited number of measurement points. We compare this scheme to standard raster scanning, as a function of the number of points, and demonstrate the advantages of the machine learning scheme.
I Results
I.1 Setup, simulations and models
Our approach relies on training a machine learning model, with real and simulated data, to enable efficient identification of the resonance frequencies in the measured signal. We employ a relevant MLP model and compare it to regular raster scanning as a function of the number of data points (subsampling).
Ninety six full raster scans were measured in an epi-illumination wide field setup (see Fig. 1a and Sec. III for details), each one under a different externally applied magnetic field which we measure through standard ODMR (since the different fields lead to different resonance frequencies).
Figure 2: In LABEL:sub@fig:nnSynth_2 and LABEL:sub@fig:nnlfIntSynth10k, ML networks (purple squares) were trained with 10 0001000010\,00010 000 full-length synthetic samples, validated on 2 00020002\,0002 000 samples and tested on other 2 00020002\,0002 000 samples to report the results. The same 2 00020002\,0002 000 test samples are used to calculate the normalized error of raster scanning (blue circles). In LABEL:sub@fig:nnSynth_2 the dataset is subsampled to the specified number of data points and a different network is trained and tested for each subsampling ratio. Accordingly, the raster scan was calculated directly on the subsampled data. In LABEL:sub@fig:nnlfIntSynth10k only one network is trained with the $600$ data points of the full dataset, then the subsampled data is linearly interpolated to upsample it to the full size for the test. Accordingly, the same procedure is used for the raster scan. LABEL:sub@fig:realdata Comparison between error of neural networks and normalized error of raster scans for different number of data points. The networks (purple squares) were trained with subsampled datasets, like as the the ones in LABEL:sub@fig:nnSynth_2, however, the training dataset contained 1 00010001\,0001 000 synthetic samples and 50 real data samples, validation was done with 46 real data samples. The same 46 raster were subsampled, the averaged error for those is depicted in blue.
Synthetic data was generated based on a simplified NV Hamiltonian [1] considering the Zeeman shift: magnetic fields at various angles ($\theta$ and $\phi$, respectively longitudinal and azimuthal, with respect to the diamond surface), were projected on the four NV orientations, determining the resonance frequencies. Resonance widths, contrast, and Gaussian noise were chosen to mimic the lineshape of the measured data, producing full ESR spectra with six or eight Lorentzians, similar to the one depicted in Fig. 1d.
The neural network’s input layer dimension ($\lambda_{n}$ in  Fig. 1e) is determined by the number of data points in an ESR spectrum and therefore was changed as networks were trained for different subsamplings: starting from the full 600-point spectrum, we subsample by taking every other point, every third point and so on, obtaining at the end spectra with 300 points, 200 points and so on. The output layer always consists of eight values that predict the central frequencies of the eight Lorentzians ($y\;^{\prime}_{i}$). These predictions are then compared to the values extracted from the full ESR spectrum, which is defined to be the ground truth, and the error is defined to be the averaged absolute value of the differences.
To accomplish a fair comparison, a similar process was performed for raster scans: the resonances extracted from subsampled data (through standard Lorentzian curve fitting) were compared to the ones extracted from the full spectrum, and the error is defined in the same way. It is important to note that the subsampled datasets were compared to the full data from which they were derived, and the error in the full scan (with the maximal number of points) is the fit error.
Some of the simulated data samples had overlapping or partly overlapping Lorentzians, for which the centers were sometimes too close to be distinguished. Note that for these data samples the neural network still outputs eight values. However, trying to identify eight resonances using a raster scan, especially with lower SNR, would often fail. Raster scans also fail when heavily subsampled, as one or more of the Lorentzians are no longer recognizable. To account for this statistical trait of the raster scans, we normalize the error by the square root of the success probability, essentially following the definition of the sensitivity and considering the success probability in terms of time (translating into additional measurements needed, as further explained in Sec. III).
I.2 Scaling and comparison
Fig. 2a depicts the ML error (purple squares) for networks that were trained with 10 0001000010\,00010 000 samples and validated with 2 00020002\,0002 000 samples to find the optimal hyperparameters (see Sec. III.2 for details on the model and its hyperparameters). As previously described, here each network was trained on the relevant subsampled data. Blue circles depict the normalized error of subsampled raster scanning, averaged over the same 2 00020002\,0002 000 samples that were used for the network’s validation. The raster scan error is distinctly high, due to the above-mentioned normalization of the error with the success probability, since in some of the data samples it was not possible to identify 8 Lorentzians. The ML model exhibits better results compared to raster scanning with an error that is more than 400 KHz better (except for the case of full length data), and remarkably, with only $10\%$ of the data points, the ML error is still below $1$ MHz. In addition, the ML error has better scaling as a function of the number of data points.
In Fig. 2b we present an alternative training scheme: in this case, just one network was trained and only on a full length dataset (600 frequency points) with 10 0001000010\,00010 000 samples and validated with 2 00020002\,0002 000 samples. To comply with the structure of the MLP, the network was tested on 2 00020002\,0002 000 subsampled dataset, that were first linearly interpolated back to full length and, only then, introduced to the network. The averaged error of these is presented in the plot (purple squares). The same 2 00020002\,0002 000 samples were used to test raster scanning (blue circles), these were also sub-sampled, interpolated and then assessed. The interpolation surprisingly improves the error for raster scanning. In fact, even though it introduces additional noise to the data, the fitting works significantly better, and for a higher number of points, the raster’s error and standard deviation are very similar to the network. However, below 120 data points the ML models shows better results compared to the raster scan. Interpolating data saves on training time, since one network fits all datasets, but at the cost of less favourable performance. In fact, when a model is trained on the full 600-points data, it learns patterns that can be distorted with the subsampling and the subsequent interpolation. For high numbers of points, the two methods are almost equivalent. Although, for lower numbers of points, starting at 120 points, a network trained for the specific number of data points gives a lower averaged error and standard deviation, as can be seen from the comparison of the two graphs in Fig. 2a (and depicted also in Fig. S1b in the supplementary material).
We note that, while not realized here, a hybrid approach might be considered and could be beneficial, wherein several networks could be trained on different subsampling sizes but not on all of them. In this case, the data to be analyzed could be interpolated to the closest available dimension and the corresponding trained network can be employed for the prediction. This can save training time while keeping the predictions less affected by the effects of the interpolation on the data.
Fig. 2c depicts the result of a network (purple squares) tested on real data, and the subsampling of real raster scans. The networks were trained (as in Fig. 2a) for specific subsamplings, with a combination of 1 00010001\,0001 000 synthetic data samples along with 50 real data samples, and was tested on 46 real data samples. The same 46 real samples were used to calculate the raster’s normalized error. The raster’s results in Fig. 2c, are better than the previous two, this is mostly due to the fact that in all 46 data samples, the eight Lorentzians were well separated. The network’s error here still has a better scaling than the raster’s error, however, for a high number of points the network error is about $1$ MHz worse than the raster, which is in agreement with training predictions that are further explained in the following section.
I.3 Analysis and generalization
We now turn to a more detailed analysis ML model behavior, further examining the behaviour of the network with regards to other parameters: the size of the training dataset, noise level and Lorentzian widths, which can vary between diamond samples due to coherence properties of the NVs [6].
Fig. 3 depicts the evolution of the validation error during training of the neural network for two dataset sizes: in green, the network was trained with 10 0001000010\,00010 000 simulated samples and validated with 2 00020002\,0002 000 samples. In red, the network was trained with 1 00010001\,0001 000 simulated samples and validated with 200 simulated samples. In blue, 50 real data samples were added to the training set, and the validation was performed on 46 real data samples. Even though training with 1 00010001\,0001 000 samples is not sufficient, as the minimal error is around $2$ MHz, there is agreement between the results on synthetic data and the ones on real data. This plot explains the difference between results depicted in Fig. 2c and Fig. 2a - respectively represented here by the blue and green curves, with a difference of $0.5-2$ MHz between them.
Figure 3: Evolution of error during the training of ML models, calculated on the validation sets of different datasets. In green (Valid. 10k): training with 10 0001000010\,00010 000 synthetic samples and validated on 2 00020002\,0002 000 synthetic samples. In red (Valid. 1k): training with 1 00010001\,0001 000 synthetic samples and validated on 200 synthetic samples. In blue (Valid. 1k real): training with 1 00010001\,0001 000 synthetic samples plus 50 real samples, validated on 46 different real samples.
Two important parameters which characterize the resonances measured in ESR experiments are the SNR and the Lorentzian widths. These vary between experimental setups, diamond samples and even individual measurements due to different noise sources. Fig. 4 presents the average error of two networks: in Fig. 4a the network was trained on a simulated dataset in which all 10 0001000010\,00010 000 samples had a width of 10 MHz and subsequently tested on three datasets with $100$ data samples, each with different resonance widths: 6 MHz (blue), 10 MHz (green), and 15 MHz (red). When tested on $10$ MHz, the network’s averaged error is 0.5 MHz lower than it is for 6 and 15 MHz. In Fig. 4b the network was trained on a different dataset of 10 0001000010\,00010 000 samples with widths within a range: from 5 MHz to 16 MHz, and then, tested on the same three datasets of Fig. 4a. This way of training proved to be a more robust model, the errors for all the tests are comparable and the network is no longer sensitive to changes in width. Similar behaviour was observed for SNR, Fig. 4c depicts a network that was trained on data samples with SNR of 4, and tested on different samples with SNR 2.5, 4 and 10. Fig. 4d depicts the test results of a network that was trained on samples with SNR randomly sampled within the range [2.5, 10]. Again, the network that was trained on a range is more robust, and when training specifically, the variations show a higher error. The results in Fig. 4 report the averages on $100$ test samples, for different number of measurements, using the interpolation training scheme.
Figure 4: ML averaged error on simulated data for different number of measurements. All the models are trained on 10 0001000010\,00010 000 samples with $600$ data points and tested on different datasets of $100$ samples subsampled and successively interpolated to the full dimension. In particular, in LABEL:sub@fig:10mhz the network is trained on samples having 10 MHz Lorentzian width and tested on three different datasets each one having different widths: 6 MHz (blue), 10 MHz (green), and 15 MHz (red). In LABEL:sub@fig:widthrange the network is trained on samples with random width within the range $[5,16]$ MHz and tested on the same datasets of LABEL:sub@fig:10mhz. LABEL:sub@fig:noiseVarFix depicts the averaged error for a model trained on samples with SNR value of 4 and tested on different samples with SNR values of 2.5 (red), 4 (green), and 10 (blue). In LABEL:sub@fig:noiseVar, the network was trained on samples with random SNR within a range of $[2.5,10]$ and tested on the same datasets as in LABEL:sub@fig:noiseVarFix.
Changes in noise and width also have an effect on subsampling raster scans, and are described in detail in Sec. III.
II Discussion
In this paper we employ ML models to achieve high bandwidth measurements without compromising sensitivity in ESR measurements with NV centers. The neural networks exhibit an advantage over raster scans, it maintains the same error for down to fifth of the data points, while for that sub sampling rate, the raster’s error increases by about 800 KHz. The ML models still perform well with an error of less than $1$ MHz using only $10\%$ of data points. Moreover, it shows an impressive ability to predict resonance locations in the presence of overlapping Lorentzians. We show the gain and flexibility of training networks in different ways: interpolating can save on training time while training multiple networks yields a lower error and standard deviation. Furthermore, we show that, for a network to be robust, it needs to be trained on a big enough dataset that is comprised of data in a range of noise and linewidth.
In the regime of sufficient training, we find it useful to have a simulation that well-describes the physics of the NV system, such that the resulting synthetic data is comparable to measured data.
The machine learning techniques applied to this quantum setting prove to be efficient, achieving an improved trade-off between high sensitivity and high dynamic range. They can be adaptively applied to measurements to achieve the desired result. E.g., reducing the measurement time by a factor 5, while maintaining the same error, improves the sensitivity by a factor $\sim 2$; alternatively, the sensitivity can remain constant while reducing the measurement time even further. Such capabilities could have signifincant impact on a broad range of measurement scenarios with large, time-varying signals, such as characterization of circuit performance, identifying transient biological signals, and more.
III Methods
III.1 Experimental setup and data acquisition

### v2

# thanks: Corresponding author

@@ Abstract
Recent years have seen significant growth of quantum technologies, and specifically quantum sensing, both in terms of the capabilities of advanced platforms and their applications. One of the leading platforms in this context is nitrogen-vacancy (NV) color centers in diamond, providing versatile, high-sensitivity, and high-spatial-resolution magnetic sensing. Nevertheless, current schemes for spin resonance magnetic sensing (as applied by NV quantum sensing) suffer from tradeoffs associated with sensitivity, dynamic range, and bandwidth. Here we address this issue, and implement machine learning tools to enhance NV magnetic sensing in terms of the sensitivity/bandwidth tradeoff in large dynamic range scenarios. Our results indicate a potential reduction of required data points by at least a factor of 3, while maintaining the current error level. Our results promote quantum machine learning protocols for sensing applications towards more feasible and efficient quantum technologies.
Quantum Sensing, Magnetic Sensing, Machine Learning, Quantum Machine Learning, Nitrogen Vacancy (NV) centers, Neural Networks
I Introduction
@@
In the context of quantum sensing, one of the leading systems is based on Nitrogen Vacancy (NV) color centers in diamond [1], which provide a versatile platform for diverse quantum sensing, notably magnetic sensing [2]. NVs have found important applications in magnetic sensing, covering paleomagnetometry [3], biosensing [4] nuclear magnetic resonance [5] and more.
Quantum sensing with NVs is realized through the illustrated spin resonance measurement techniques. Although this approach achieves quantitative vectorial information with high sensitivity and spatial resolution, it suffers from a trade-off between sensitivity and bandwidth, specifically in the high-dynamic range regime. In fact, working with small fields (small-dynamic range) enables an optimal sensing strategy, which relies on precise measurements at a predetermined high-sensitivity point (the point of maximal signal gradient). However, this is not possible in the regime of large-dynamic range signals. As shown in Fig. 1, the standard approach to quantitatively measure vectorial magnetic fields using NVs is essentially a common spin resonance measurement termed Electron Spin Resonance (ESR), usually performed through optical readout for NVs, referred to as Optically Detected Magnetic Resonance (ODMR), using an experimental setup schematically depicted in Fig. 1a [6].
@@
Figure 1: LABEL:sub@fig:scheme-setup Experimental setup schematic; MW pulses are used to identify the spin resonances of the NV center within the diamond sample. The resonance frequencies are related to the external magnetic field orientation and intensity through the Zeeman shift. Green excitation laser illuminates the diamond and fluorescence from the NV centers is collected and read by a detector. LABEL:sub@fig:scheme-nv The NV can appear in 4 possible orientations within the diamond lattice. LABEL:sub@fig:scheme-level Schematic of the NV’s energy levels. Under an external magnetic field, the degeneracy between the $m_{s}=\pm 1$ energy levels of the ground state is lifted. MW pulses drive the transitions between $m_{s}=0$ and $m_{s}=\pm 1$. The ground state is excited with a green laser (532 nm) and can decay via two pathways: one is non-radiative and non-spin-conserving through the singlet state, while the other is spin-conserving and emits red fluorescence in the range of 650-800 nm. The $m_{s}=0$ state decays mostly radiatively, while $m_{s}=\pm 1$ includes a significant non-radiative decay through the singlet state into the ground spin state $m_{s}=0$. LABEL:sub@fig:scheme-esr ESR spectrum, with measured data in blue and the corresponding fit in purple (to a set of 8 Lorentzian functions). In black, a subset of the data, i.e $\lambda_{i}$. The resonance frequencies are marked with $y_{i}$. LABEL:sub@fig:scheme-net Schematic of the neural network, the inputs $\lambda_{i}$ are the measured data points and the outputs are the predictions $y^{\prime}_{i}$ of the 8 resonance frequencies $y_{i}$.
@@
The orientation of NV defects within the crystal lattice (Fig. 1b) and its energy level structure (Fig. 1c), enable the optical detection of NV spin resonances, realized through the combined application of green-light excitation and microwave (MW) irradiation, while detecting the NV spin-dependent red fluorescence. As depicted in Fig. 1d, a full ESR spectrum of an NV ensemble consists of 8 resonances, representing 2 resonant frequencies for each of the 4 crystallographic orientations the NV can take in a single crystal diamond sample (Fig. 1b). Basically, this measurement amounts to identifying resonance positions in the frequency space. Once obtained, a full vectorial magnetic field can be extracted. Further details regarding the NV system, the experimental setup (Figs. 1a and 1c) and the ODMR measurement scheme can be found in Sec. II.
@@
Identifying the resonance positions requires scanning the signal over the frequency space, often referred to as raster scanning. This scan must cover the entire relevant frequency range, determined by the desired dynamic range. At higher fields, the Zeeman splitting increases, resulting in larger separations between resonances. This enables a wider dynamic range, allowing the system to measure both small and large magnetic fields without overlapping resonances. The scan resolution defined by the number of data points within the frequency range, depends on the resonance widths and Signal to Noise Ratio (SNR). There is a minimal scan resolution required to successfully retrieve the resonance frequencies, and in general the measurement error depends on these various parameters (linewidth, SNR, scan resolution) as detailed in Sec. II.
@@
The sensitivity for an ODMR measurement is defined to be $\eta=(\delta B*\sqrt{T})$ ; or equivalently in terms of frequency, $\eta=(\delta\nu*\sqrt{T})/\gamma$, where $\delta B$ is the uncertainty in magnetic field $\delta\nu$ is the error in frequency, T is the measurement time which corresponds to the number of data points, averaging and SNR, and $\gamma$ is the gyromagnetic ratio of the NV [7]. A shorter measurement time, while maintaining the same SNR, improves sensitivity and enables the measurement of time-varying signals. Reducing measurement time could be achieved by improving the experimental setup, i.e., having higher SNR or better NV properties. However, it is not less important to consider the optimal number of data points in a measurement.
In this paper, we propose an Machine Learning (ML) algorithm as a promising approach to address the current trade-off. We demonstrate that training an appropriate neural network using a combination of real and simulated data enables an improvement in measurement bandwidth for a given sensitivity goal, in the large dynamic range scenario.
ML is a research field that deals with the development of artificial intelligence methods that learn from data. Several approaches were developed to solve the two main supervised ML tasks of classification, i.e. to predict a categorical value associated to some predictors (called features in the ML jargon), and regression where the predicted value is continuous [8].Our work lies within the emerging field of Quantum Machine Learning (QML), positioned at the intersection of ML quantum physics and quantum computing [9]. In general, QML encompasses various approaches, including the application of classical ML models to quantum systems, for example, in noise analysis of quantum devices [10], or the implementation of ML models on quantum devices, for example, quantum generative models [11].
Neural networks are powerful ML models constituted by interconnected layers of artificial neurons that are trained to minimize a loss function [12]. Recently, thanks to the advances in computational power and the growing availability of large datasets, neural networks have outperformed other ML algorithms, becoming a general-purpose tool for either clustering, classification, regression [13] and generative [14] tasks. In fact they are capable of universal function approximation [15] and easily adapt to different scenarios. Their flexibility allows them to serve as the underlying structure in various architectures and tasks. In our work we adopt a classical feed-forward architecture, commonly known as Multilayer Perceptron (MLP[16], as illustrated in Fig. 1e. The network is trained using gradient descent to minimize a loss function between the prediction and the desired outputs. More details can be found in Sec. II.2. Such models have already been used with NV centers for sensing [17], and for noise spectroscopy [18]. These approaches hold promise for various practical applications, such as in the study of integrated circuits [19] and medical diagnostics [20]. In our work, we apply neural networks to the full reconstruction of an external magnetic field using an NV ensemble. We analyze the network’s performance for a decreasing number of measurement points, while exploring different training dataset sizes, and altering noise and lineshape conditions. We compare these results to similar analysis done for raster scans and find that the network has advantages over raster scanning. With the right training, the network can be insensitive to variations in noise and lineshape. Moreover, the network’s measurement error is more robust and scales better for fewer data points.
II Methods
II.1 Experimental setup and data acquisition